OpenStudy (anonymous):

Given f(x) = x + 2 and g(x) = x2 – 4, find the function (g/f)(x).

OpenStudy (jhannybean):

\[\large (\frac{g}{f})(x)=\frac{x^2-4}{x+2}\]

OpenStudy (jhannybean):

expand the numerator. \[\large \frac{(x-2)(x+2)}{(x+2)}\] now eliminate like-terms. \[\large \frac{(x-2)\cancel{(x+2)}}{\cancel{(x+2)}} \] therefore you get \[\large (\frac{f(x)}{g(x)})(x) = x-2\]

OpenStudy (jhannybean):

See how that works?

OpenStudy (anonymous):

Yeah, would the excluded value be -2 as well? or just 2? @Jhannybean

OpenStudy (anonymous):

I'm guessing it's just 2, since that was what your crossed out. Or what was excluded.

OpenStudy (jhannybean):

Umm..what is excluded value again ><

OpenStudy (anonymous):

idk, it asks me for the excluded value as well. :P

OpenStudy (jhannybean):

Ohh, the values excluded in domain ? that would mean x cannot = - 2,i believe.

OpenStudy (anonymous):

ohhh, okay thanks!

OpenStudy (jhannybean):

And the root is x=2. I believe you're right!

OpenStudy (anonymous):

cool thanks!